# Chaos in Systems with Quantum Group Symmetry

### v1

@@ 1 Introduction
Quantum group (QG) is a generalization of the usual groups which allows the commutation relation of the algebra elements to be non-linear. Initially quantum groups were discovered in connection with integrability, [1, 2], see [3, 4] for reviews. To the best of our knowledge, all physical models in which the action of a QG was studied were integrable. Recently in [5, 6] it was emphasized that certain theories have global symmetries given by a QG, meaning that its charges commute with a Hamiltonian and states and operators fall into irreducible representations. In this case the number and type of constraints that QG imposes on the observables of a theory is similar to that of a usual global symmetry and the number of integrals of motion is finite. This suggests that there should be models with QG symmetry that are not integrable. Nevertheless, the particular theory considered in [5, 6] as a main example is also integrable. In this paper we construct a model that is QG-invariant and manifestly chaotic. We do so by starting with QG-invariant version of the XXZ spin chain studied in [7] and deform it by an operator which preserves QG $U_{q}(sl_{2})$ but breaks integrability.
@@
It is now timely to mention that the theory we study in this paper is not unitary and the Hamiltonian is not Hermitian. Our choice is motivated by connection to the model studied in [5, 6] which is known to have a continuum limit when $q$ is on a unit circle. Instead of unitarity our model possesses the so-called PT-symmetry [9]. PT-symmetric models have recently attracted an increased interest due to a connection with open quantum systems, including their experimental realizations, see [10] and references therein. This symmetry also plays an important role in Lagrangian description of strongly-coupled CFTs as most recently discussed in [11, 12]. Classification of non-unitary chaotic systems is much more complicated, and consists of as many as 38 classes [13], [14]. While non-unitary systems are certainly more exotic, they can exhibit some interesting phenomena not present in the unitary ones. For example, as a function of the coupling constant our model undergoes a transition to the so-called PT-symmetry broken phase [15], in which the eigenvalues are no longer real, but some of them come in complex-conjugate pairs. The co-existence of real and complex eigenvalues also makes the statistical properties much richer.
@@ 2 Spin chain with $U_{q}(sl_{2})$ symmetry and its algebraic structure
The formulas (8) and (7) specify the universality class of the complex Hamiltonian. [^4] The symmetry class is BDI in the notations of [14, Appendix B]. Since $\mathcal{H}_{0}$ is integrable, for any value of $q$ we expect the Poisson distribution for the level spacing.
@@ 3 Next-to-nearest deformation preserving symmetries
It is easy to check that in the limit $q\to 1$ they become exactly (11). Finally, we write the next-to-nearest deformation term as, [^6] In terms of Temperley-Lieb generators it can be equivalently written as $V(\lambda)=\lambda\sum_{i=1}^{L-2}\bigg(e_{i}e_{i+1}+e_{i+1}e_{i}-\frac{q+q^{-1}}{2}(e_{i}+e_{i+1})+\frac{q^{2}+q^{-2}}{4}\mathbb{1}\bigg)$
@@ 5.2 Analysis of different types of eigenvalues
$z_{k}=\frac{E_{i}^{\rm NN}-E_{i}}{E_{i}^{\rm NNN}-E_{i}}\equiv r_{k}e^{i\theta_{k}},$ (32)
where for a given eigenvalue $E_{i}$ one has to find its nearest neighboring eigenvalue $E_{i}^{\rm NN}$ in the complex plane as well as its next-to-nearest neighboring eigenvalue. By definition $z$ is bounded in the unit circle, $z\leq 1$. Now one can study a distribution $P(z)$ in the complex plane. The repulsion of complex eigenvalues in the chaotic regime manifests itself in the CLSR distribution as well. See for example [37] for the plots of complex spacing ratio for generic complex matrices, so called GinUE ensemble. An important quantity that we compute out of the distribution of $z_{k}$ is their average absolute value $\braket{r}$. This single number is used to distinguish between integrable and chaotic behavior. An advantage of $\braket{r}$ is that it can be computed both for the complex spectrum and for the real spectrum. [^9] Note that when computed for the real spectrum, the average value of $\braket{r}$ does not have to coincide with average real spacing ratio $\braket{\tilde{r}}$. That is, this analysis is applicable for all values of $q$ and $\lambda$. The value of $\braket{r}$ for different ensembles is presented in the Table 1.
@@ Acknowledgements
We thank Barak Gabai, Andreas Läuchli, Jiaxin Qiao, Bernardo Zan for discussions. AZ would like to thank Elena Petrova for valuable discussions and for suggesting useful references. The work of VG is supported by Simons Foundation grant 994310 (Simons Collaboration on Confinement and QCD Strings).

### v2

@@ 1 Introduction
Quantum group (QG) is a generalization of the usual groups which allows the commutation relation of the algebra elements to be non-linear. Initially quantum groups were discovered in connection with integrability, [1, 2], see [3, 4] for reviews. To the best of our knowledge, all physical models in which the action of a QG was studied were integrable, even though some systems with such symmetry have a high chance of having chaotic behavior [5]. Recently in [6, 7] it was emphasized that certain theories have global symmetries given by a QG, meaning that its charges commute with a Hamiltonian and states and operators fall into irreducible representations. In this case the number and type of constraints that QG imposes on the observables of a theory is similar to that of a usual global symmetry and the number of integrals of motion is finite. This suggests that there should be models with QG symmetry that are not integrable. Nevertheless, the particular theory considered in [6, 7] as a main example is also integrable. In this paper we construct a model that is QG-invariant and manifestly chaotic. We do so by starting with QG-invariant version of the XXZ spin chain studied in [8] and deform it by an operator which preserves QG $U_{q}(sl_{2})$ but breaks integrability.
@@
It is now timely to mention that the theory we study in this paper is not unitary and the Hamiltonian is not Hermitian. Our choice is motivated by connection to the model studied in [6, 7] which is known to have a continuum limit when $q$ is on a unit circle. Instead of unitarity our model possesses the so-called PT-symmetry [10]. PT-symmetric models have recently attracted an increased interest due to a connection with open quantum systems, including their experimental realizations, see [11] and references therein. This symmetry also plays an important role in Lagrangian description of strongly-coupled CFTs as most recently discussed in [12, 13, 14]. Classification of non-unitary chaotic systems is much more complicated, and consists of as many as 38 classes [15], [16]. While non-unitary systems are certainly more exotic, they can exhibit some interesting phenomena not present in the unitary ones. For example, as a function of the coupling constant our model undergoes a transition to the so-called PT-symmetry broken phase [17], in which the eigenvalues are no longer real, but some of them come in complex-conjugate pairs. The co-existence of real and complex eigenvalues also makes the statistical properties much richer.
@@ 2 Spin chain with $U_{q}(sl_{2})$ symmetry and its algebraic structure
The formulas (8) and (7) specify the universality class of the complex Hamiltonian. [^4] The symmetry class is BDI in the notations of [16, Appendix B]. Since $\mathcal{H}_{0}$ is integrable, for any value of $q$ we expect the Poisson distribution for the level spacing.
@@ 3 Next-to-nearest deformation preserving symmetries
It is easy to check that in the limit $q\to 1$ they become exactly (11). Finally, we write the next-to-nearest deformation term as, [^6] In terms of Temperley-Lieb generators it can be equivalently written as $V(\lambda)=\lambda\sum_{i=1}^{L-2}\bigg(e_{i}e_{i+1}+e_{i+1}e_{i}-\frac{q+q^{-1}}{2}(e_{i}+e_{i+1})+\frac{q^{2}+q^{-2}}{4}\mathbb{1}\bigg)$ , [^7] A certain different choice of coefficients might yield an integrable deformation. Higher spin conserved charge quadratic in Temperley-Lieb generators was considered in [33].
@@ 5.2 Analysis of different types of eigenvalues
$z_{i}=\frac{E_{i}^{\rm NN}-E_{i}}{E_{i}^{\rm NNN}-E_{i}}\equiv r_{i}e^{i\theta_{i}},$ (32)
where for a given eigenvalue $E_{i}$ one has to find its nearest neighboring eigenvalue $E_{i}^{\rm NN}$ in the complex plane as well as its next-to-nearest neighboring eigenvalue. By definition $z$ is bounded in the unit circle, $|z|\leq 1$. Now one can study a distribution $P(z)$ in the complex plane. The repulsion of complex eigenvalues in the chaotic regime manifests itself in the CLSR distribution as well. See for example [40] for the plots of complex spacing ratio for generic complex matrices, so called GinUE ensemble. An important quantity that we compute out of the distribution of $z_{i}$ is their average absolute value $\braket{r}$. This single number is used to distinguish between integrable and chaotic behavior. An advantage of $\braket{r}$ is that it can be computed both for the complex spectrum and for the real spectrum. [^10] Note that when computed for the real spectrum, the average value of $\braket{r}$ does not have to coincide with average real spacing ratio $\braket{\tilde{r}}$. That is, this analysis is applicable for all values of $q$ and $\lambda$. The value of $\braket{r}$ for different ensembles is presented in the Table 1.
@@ 5.3.2 Complex spectrum
Some of the CLSR plots also exhibit a “bow and arrow” feature: an increased density along an “arc”, a straight vertical line at $\text{Re}\;z=\frac{1}{2}$, and the real line. This feature is explained by PT-symmetry [42, Appendix D], and was also observed in [43]. The values of $z_{k}$ start to form an arc when two eigenvalues coincide and move into a complex plane in the opposite directions, while having the third nearest eigenvalue on the real line. The values of $z_{k}$ transition from the arc to the vertical line when two complex conjugated eigenvalues start being further away from each other than from the third real one. We make more plots for different values of $q$ to investigate this structure in Appendix B.
@@ Acknowledgements
We thank Barak Gabai, Igor Klebanov, Andreas Läuchli, Rafael Nepomechie, Bendeguz Offertaler, Jiaxin Qiao, Ruth Shir, Grigory Tarnopolsky, Bernardo Zan for discussions and comments on the draft. AZ would like to thank Elena Petrova for valuable discussions and for suggesting useful references. The work of VG is supported by Simons Foundation grant 994310 (Simons Collaboration on Confinement and QCD Strings).
@@ Appendix A IR phase of the spin chain
Appendix B Complex level spacing ratio distributions
In this appendix we present more numerical results of complex level spacing ratio (CLSR) distributions $P(z)$ for different values of parameter $q$ and for different values of coupling $\lambda$, see Figure 8. Part (a) of the figure shows CLSR for $q=e^{i\pi/6.7}$; part (b) shows CLSR for $q=e^{i\pi/2.7}$, this is in addition to $q=e^{i\pi/4.8}$ shown in the main text. The difference between these values of the parameter $q$ is the number of eigenvalues that become complex with the increase of $\lambda$. In (a) the maximal number of complex eigenvalues reaches 335 out of 1688, that is, the majority of eigenvalues stays real for any value of $\lambda$, see Table 2. In (b) this number is 1544 out of 1688, which makes almost all eigenvalues to be complex in the most chaotic regime. Moreover, in (b) the PT-symmetry breaking happens for very small values of $\lambda$, which means that for all plots presented in (b) the majority of eigenvalues is complex.
Despite having different values of $q$ and therefore different number of complex eigenvalues, both (a) and (b) exhibit the “bow and arrow” structure, an artifact of PT symmetry (see discussion around Figure 4). Such a structure is best visible both for small and very large values of the deformation parameter $\lambda$. This structure emerges when there are enough triples of eigenvalues close to each other, one real and two complex conjugated ones. Interestingly, this structure disappears or becomes indistinguishable from the rest of the distribution for the most chaotic regime, where the values of $\lambda$ are between $0.5$ and $5$. It would be interesting to investigate further any correlation between “bow and arrow” structure and integrability-chaos transition.
Figure 8: Complex level spacing ratio distributions $P(z)$ for different values of $\lambda$. Parameter $q$ is: (a) $q=e^{i\pi/6.7}$; (b) $q=e^{i\pi/2.7}$. Spin chain length is $L=16$. Eigenvalues are from $q$-parity even sector with $\{\ell=1,S^{z}=1\}$.